# Instruction Pre-Training: Language Models are Supervised Multitask Learners

### v1

@@ Abstract
Daixuan Cheng  Yuxian Gu  Shaohan Huang🖂  Junyu Bi  Minlie Huang🖂  Furu Wei Microsoft Research               Tsinghua University https://huggingface.co/instruction-pretrain
footnotetext: 🖂 Corresponding Author.
@@ 2.2 LM Pre-Training
After collecting the synthesized instruction-response pairs, we employ templates from Longpre et al. (2023) to diversify instruction formats, and templates from Cheng et al. (2023) to concatenate each raw text with its instruction-response pairs. As shown in Figure 3, by concatenating the texts and instruction-pairs from $M$ rounds, we create an $M$-shot example for subsequent pre-training.
Except for the pre-training data, Instruction Pre-Training keeps all other pre-training settings the same as Vanilla Pre-Training: training with the next-token prediction objective (Radford et al., 2018) and computing loss on all tokens. We conduct both general pre-training from scratch and domain-adaptive continued pre-training to verify the effectiveness in different pre-training scenarios.
@@
Table 1: General Performance of the Pre-Trained Base Models via Vanilla Pre-Training (Vanilla PT), mixing raw corpora with fine-tuning data for the instruction synthesizer (Mix PT), and Instruction Pre-Training (Instruct PT) in general pre-training from scratch.
@@ 3.2 General Pre-Training From Scratch
To create instruction-augmented corpora, we conduct two rounds of instruction synthesis, converting 1/5 of the raw corpora (40M raw texts) into instruction-augmented texts. The first round converts 20M raw texts, and the second round uses the raw texts and instruction-response pairs from the first round to convert another 20M raw texts. The resulted corpora contain 200M synthesized pairs. An example of a 2-shot instruction-augmented text is shown in Table 12 in Appendix.
@@ 3.3 Domain-Adaptive Continual Pre-Training
We conduct 3-round inference to covert all the domain-specific corpora. Each round processes 1/3 of the raw corpora, inheriting the raw texts and instruction-response pairs from previous rounds. Examples of the instruction-augmented texts are in Table 13 and 14 in Appendix.
@@
We continue to pre-train Llama3-8B on each domain respectively, detailed settings are in Table 10 in Appendix. We follow the same setting in Cheng et al. (2023) to evaluate models on the domain-specific tasks. Detailed evaluation settings are in Appendix C.
@@ 4.1 General Pre-Training From Scratch
Vanilla PT 62.9 84.7 82.2 65.4 64.9 72.0
Instruct PT 74.6 87.1 82.4 65.7 63.6 74.7
Table 3: Domain-Specific Task Performance of Llama3-8B without continued pre-training, after continued pre-training via Vanilla Pre-Training (Vanilla PT), and after continued pre-training via Instruction Pre-Training (Instruct PT). Both Vanilla PT and Instruct PT mix domain-specific corpora with general instructions to boost prompting ability. We also display the performance of Llama3-70B for reference.
@@ 5.2 Instruction-Augmented Corpora
As shown in Table 6, our instruction synthesizer generates instruction-response pairs spanning 49 different task categories, with over 85% relevance to the context and 70% response accuracy. We further group the task categories into 9 general task scenarios. Figure 6 shows the percentages of each task scenario in the instruction augmented corpora for general pre-training. Our synthesized tasks cover all general task scenarios, demonstrating the effectiveness of our instruction synthesizer in generating a highly diverse tasks.
@@ 6 Related Work
Data curation for LM pre-training typically involves collection, cleaning, and organization. Most pre-training data are collected from the Internet to ensure diversity (Raffel et al., 2020; Penedo et al., 2023; Wenzek et al., 2020; Gao et al., 2020). Although diverse, web-scraped data often contain low-quality and duplicate content. Therefore, data cleaning techniques are applied to these corpora, including language identification (Joulin et al., 2016), perplexity-based (Wenzek et al., 2020), classifier-based (Brown et al., 2020), and rule-based (Raffel et al., 2020; Rae et al., 2021) filtering. Data organization aims at performing more fine-grained programming of the data, including data selection Albalak et al. (2024); Xie et al. (2024) and constructing training instances related to downstream usage Gu et al. (2022a, 2023); Shi et al. (2023); Jiang et al. (2024b). Our work explores an orthogonal direction: augmenting raw corpora with large-scale supervised signals.
@@ 7 Conclusion
This paper proposes Instruction Pre-Training to explore supervised multitask learning for pre-training. Instead of directly pre-training on raw corpora, Instruction Pre-Training augments the corpora with instruction-response pairs generated by an instruction synthesizer, then pre-trains LMs on the augmented corpora. Our instruction synthesizer, fine-tuned from a highly diverse data collection, is capable of generating diverse instruction-response pairs from various corpora. In general pre-training from scratch, Instruction Pre-Training not only outperforms Vanilla Pre-Training on the pre-trained base models but also benefits more from further instruction tuning. In continual pre-training, Instruction Pre-Training substantially enhances the performance of Llama3-8B in two different domains. Looking ahead, we hope our work can inspire further exploration into this promising area of supervised multitask pre-training, effectively enhancing the general abilities of LMs.
@@ Appendix B Tuning and Inference Settings for Instruction Synthesizer
Let’s think step by step. $<$ANS$>r$CoT}
@@
During each round of inference, we concatenate the formatted examples from previous rounds with the formatted raw text of the current round as the input for the instruction synthesizer. Subsequently, the instruction synthesizer generates a sequence of instruction-response pairs. The maximum sequence length for inference corresponds to that of the target LM intended for pre-training. We use the vLLM (Kwon et al., 2023) framework for acceleration. It takes about 2 days to synthesize instruction-response pairs for 1B tokens of raw corpora on a single A100-80GB GPU.
Hyperparameter Assignment
@@ Appendix C LM Evaluation
We follow the same prompting settings of AdaptLLM (Cheng et al., 2023): for biomedicine domain, we evaluate zero-shot performance on PubMedQA (Jin et al., 2019) and USMLE (Jin et al., 2021), few-shot performance on ChemProt (Kringelum et al., 2016), MQP (McCreery et al., 2020) and RCT (Dernoncourt and Lee, 2017); for finance domain, we evaluate few-shot performance on ConvFinQA (Chen et al., 2022), FPB (Malo et al., 2014), FiQA SA (Maia et al., 2018), Headline (Sinha and Khandait, 2021), and NER (Alvarado et al., 2015).

### v2

@@ Abstract
Daixuan Chengthanks:  Contribution during an internship at the CoAI Group, Tsinghua University. 🖂 Corresponding Author. Yuxian Gu Shaohan Huang🖂  Junyu Bi Minlie Huang🖂 Furu Wei Microsoft Research The CoAI Group, DCST, Institute for Artificial Intelligence, State Key Lab of Intelligent Technology and Systems, Beijing National Research Center for Information Science and Technology, Tsinghua University daixuancheng6@gmail.com    guyx21@mails.tsinghua.edu.cn {shaohanh,fuwei}@microsoft.com   aihuang@tsinghua.edu.cn https://huggingface.co/instruction-pretrain
@@ 2.2 LM Pre-Training
After collecting the synthesized instruction-response pairs, we employ templates from Longpre et al. (2023) to diversify instruction formats, and templates from Cheng et al. (2023) to concatenate each raw text with its instruction-response pairs. As shown in Figure 3, by concatenating the texts and instruction-response pairs from $M$ rounds, we create an $M$-shot example for subsequent pre-training.
Except for the pre-training data, Instruction Pre-Training keeps all other pre-training settings the same as Vanilla Pre-Training: training with the next-token prediction objective (Radford et al., 2018) and computing loss on all tokens. We conduct both general pre-training from scratch and domain-adaptive continual pre-training to verify the effectiveness in different pre-training scenarios.
@@
Table 1: General performance of the pre-trained base models via Vanilla Pre-Training (Vanilla PT), mixing raw corpora with fine-tuning data for the instruction synthesizer (Mix PT), and Instruction Pre-Training (Instruct PT) in general pre-training from scratch. All the pre-training methods use the same number of tokens for model training.
@@ 3.2 General Pre-training From Scratch
To create instruction-augmented corpora, we conduct two rounds of instruction synthesis, converting 1/5 of the raw corpora (40M raw texts) into instruction-augmented texts. The first round converts 20M raw texts, and the second round uses the raw texts and instruction-response pairs from the first round to convert another 20M raw texts. The resulted corpora contain 200M synthesized pairs amounting to about 10B tokens. An example of a 2-shot instruction-augmented text is shown in Table 16 in Appendix.
@@ 3.3 Domain-Adaptive Continual Pre-Training
We conduct 3-round inference to covert all the domain-specific corpora. Each round processes 1/3 of the raw corpora, inheriting the raw texts and instruction-response pairs from previous rounds. Examples of the instruction-augmented texts are in Table 17 and 18 in Appendix.
@@
We continue to pre-train Llama3-8B on each domain respectively, detailed settings are in Table 14 in Appendix. We follow the prompting settings in Cheng et al. (2023) to evaluate models on the domain-specific tasks. Detailed evaluation settings are in Appendix C.
@@ 4.1 General Pre-Training From Scratch
Vanilla PT-8B 62.9 84.7 82.2 65.4 64.9 72.0
Instruct PT-8B 74.6 87.1 82.4 65.7 63.6 74.7
Table 3: Domain-specific task performance of Llama3-8B without continued pre-training, after continued pre-training via Vanilla Pre-Training (Vanilla PT), and after continued pre-training via Instruction Pre-Training (Instruct PT). Both Vanilla PT and Instruct PT mix domain-specific corpora with general instructions to boost prompting ability, and use the same number of tokens for model training. The performance of Llama3-70B is displayed for reference.
@@ 5.2 Instruction-Augmented Corpora
As shown in Table 6, our instruction synthesizer generates instruction-response pairs spanning 49 different task categories, with over 85% relevance to the context and 70% response accuracy. We further group the task categories into 9 general task scenarios. Figure 6 shows the percentages of each task scenario in the instruction augmented corpora for general pre-training. Our synthesized tasks cover all general task scenarios, demonstrating the effectiveness of our instruction synthesizer in generating a highly diverse tasks. We conduct further analysis in Appendix E for human evaluation, Appendix D for data contamination, and Appendix F for domain distribution and diversity.
@@ 6 Related Work
Data curation for LM pre-training typically involves collection, cleaning, and organization. Most pre-training data are collected from the Internet to ensure diversity (Raffel et al., 2020; Penedo et al., 2023; Wenzek et al., 2020; Gao et al., 2020). Although diverse, web-scraped data often contain low-quality and duplicate content. Therefore, data cleaning techniques are applied to these corpora, including language identification (Joulin et al., 2016), perplexity-based (Wenzek et al., 2020), classifier-based (Brown et al., 2020), and rule-based (Raffel et al., 2020; Rae et al., 2021) filtering. Data organization aims at performing more fine-grained programming of the data, including data selection Albalak et al. (2024); Xie et al. (2024) and constructing training instances related to downstream usage Gu et al. (2022a, 2023); Shi et al. (2023); Jiang et al. (2024b); Maini et al. (2024). Our work explores an orthogonal direction: augmenting raw corpora with large-scale supervised signals.
@@ 7 Conclusion
This paper proposes Instruction Pre-Training to explore supervised multitask learning for pre-training. Instead of directly pre-training on raw corpora, Instruction Pre-Training augments the corpora with instruction-response pairs generated by an instruction synthesizer. Our instruction synthesizer, fine-tuned from a highly diverse data collection, is capable of generating diverse instruction-response pairs from various corpora. In pre-training from scratch, Instruction Pre-Training not only outperforms Vanilla Pre-Training on the pre-trained base models but also benefits more from further instruction tuning. In continual pre-training, Instruction Pre-Training substantially enhances the performance of Llama3-8B in two different domains. Looking ahead, we hope our work can inspire further exploration into this promising area of supervised multitask pre-training, effectively enhancing the general abilities of LMs.
@@
Acknowledgements
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604) and the NSFC key project (with No. 61936010).
@@ Appendix B Tuning and Inference Settings for Instruction Synthesizer
Let’s think step by step. <<<ANS>>> {CoT}
@@
During each round of inference, we concatenate the formatted examples from previous rounds with the formatted raw text of the current round as the input for the instruction synthesizer. Subsequently, the instruction synthesizer generates a sequence of instruction-response pairs. The maximum sequence length for inference corresponds to that of the target LM intended for pre-training. We use the vLLM (Kwon et al., 2023) framework for acceleration. It takes about 1 day to synthesize instruction-response pairs for 1B tokens of raw corpora on a single A100-80GB GPU.
Hyper-Parameter Assignment
@@ Appendix C LM Evaluation
We follow the prompting settings of AdaptLLM (Cheng et al., 2023): for biomedicine domain, we evaluate zero-shot performance on PubMedQA (Jin et al., 2019) and USMLE (Jin et al., 2021), few-shot performance on ChemProt (Kringelum et al., 2016), MQP (McCreery et al., 2020) and RCT (Dernoncourt and Lee, 2017); for finance domain, we evaluate zero-shot performance on ConvFinQA (Chen et al., 2022) and few-shot performance on FPB (Malo et al., 2014), FiQA SA (Maia et al., 2018), Headline (Sinha and Khandait, 2021), and NER (Alvarado et al., 2015).
ARC-e/c BoolQ SIQA WG PIQA OBQA HS MMLU
Total Eval Examples 2376/1172 3270 1954 1267 1838 500 10042 14042
Contam in Raw Corpora 5/3 144 0 0 3 0 4 20
Contam in Ins-Aug Corpora 5/4 144 0 0 3 0 4 22
Contam in Synthesized Pairs 0/1 0 0 0 0 0 0 2
Table 9: Data contamination analysis of raw corpora, instruction-augmented corpora and the synthesized instruction-response pairs. “WG” and “HS” represent WinoGrande and HellaSwag, respectively.
Accuracy Relevance # Category
General 75.5 87.5 51
BioMed. 81.0 97.0 21
Finance 73.5 88.0 39
Table 10: Human evaluation of response accuracy, context relevance, and number of task categories on the instruction-augmented corpora.
Appendix D Data Contamination Analysis
We measure cross-contamination between the evaluation datasets and the training data using the sub-string match method described in OpenAI (2023): an evaluated example is considered contaminated if a sub-string of it appears in the training data. Table 9 shows:
  • Total Eval Examples: The number of all evaluated examples in each dataset.
  • Contam in Raw Corpora: The number of contaminated examples in the raw corpora.
  • Contam in Ins-Aug Corpora: The number of contaminated examples in the instruction-augmented corpora, which includes the raw corpora and the synthesized instruction-response pairs.
  • Contam in Synthesized Pairs: The number of contaminated examples introduced by synthesized pairs, calculated by subtracting the number of contaminated examples in the raw corpora from those in the instruction-augmented corpora.
The results indicate the synthesized pairs introduce minimal contamination to the training data.
Coverage Coverage (multi-domain) Overlap
86.8 77.8 84.9
Table 11: Domain coverage and overlap between the raw text and the synthesized instruction-response pairs.
Encyclo Fiction Academic Trivia News Expert Social Code
Fine-tune Data 22.2 11.1 22.2 3.7 7.4 29.6 3.7 0.0
Raw Corprora 5.8 9.6 3.5 0.0 20.3 42.8 14.7 3.3
Synthesized Pairs 5.8 11.8 3.3 0.1 18.8 46.0 11.1 3.1
Table 12: Domain distribution of fine-tuning data for the instruction synthesizer, raw corpora and synthesized instruction-response pairs. “Encyclo", “Academic", “Expert" and “Social" represent Encyclopedia, Academic Tests, Expert Materials and social media domains, respectively.
Appendix E Human Evaluation on Instruction-Augmented Corpora
We conduct human evaluation to analyze the instruction-augmented corpora from the following aspects:
  • Response Accuracy: A binary score indicating whether the response is accurate based on the instruction and context, where 1 means accurate and 0 means inaccurate. We report the average score of all responses.
  • Context Relevance: A binary score indicating whether the instruction-response pair is relevant to the context of the raw text, where 1 means relevant and 0 means irrelevant. We report the average score of all instruction-response pairs.
  • # Task Category: The evaluator categorizes each instruction-response pair using a predefined list of task categories from Wang et al. (2022). We report the number of different categories of all the instruction-response pairs to show diversity.
From the results in Table 10, the synthesized instruction-response pairs span 51 different task categories, with over 85% relevance to the context and 70% response accuracy.
Appendix F Domain Distribution Analysis
We analyze domain distribution to evaluate the effectiveness of our instruction synthesizer in generating instruction-response pairs closely aligned with the domain of the given raw context.
Domain Coverage and Overlap
For each instruction-augmented text, we calculate the following scores and report the average on all the instruction-augmented texts:
  • Domain Coverage: The ratio of text domains included in the instruction domains to all text domains.
  • Domain Coverage (multi-domain only): We specifically compute domain coverage for the cases where a raw text contains multiple domains.
  • Domain Overlap: The overlap of raw text domains and instruction domains divided by the union of raw text and instruction domains.
As shown in Table 11, the synthesized instruction-response pairs cover most of the domains in the raw text, with a high domain overlap with the raw text. For the texts containing more than one domain, our instruction synthesizer generates, on average, 5 instruction-response pairs per raw text, with each pair potentially covering a different domain. According to the domain coverage (multi-domain only), when a single raw text includes multiple domains, our instruction synthesizer can effectively generate instruction-response pairs that cover most of the text domains.
Domain Distribution
We analyze domain distributions of the following sources:
  • Fine-tuning data for the instruction synthesizer.
  • Raw pre-training corpora (Penedo et al., 2023).
  • Synthesized instruction-response pairs based on the raw pre-training corpora.
As shown in Table 12, despite the domain distributions of fine-tuning data and raw corpora being very different, the synthesized pairs closely follow the domain distribution of the raw corpora.